# End-to-End Chess Recognition

### v1

@@ 4.2 Relative Object Detection Approach
5 Implementation Details
5.1 Classification
@@
For the classification approach, we employed a ResNeXt [29] model. Based on the ResNet [9] architecture, ResNeXt models constitute a powerful variant that can learn complex representations of images. Their introduced concept of “cardinality” (i.e. the number of parallel branches used in each residual block) both enables deeper architectures with reduced computation complexity and allows complex representations to be learned by aggregating the information of the multiple branches. Because of these modifications, this family of models can achieve impressive results in image classification.
@@
For our experiments, we trained the resnext101_32d variant, which uses a cardinality of 32 and a width (i.e. number of filters) of 8. This means that each residual block in the network contains 32 parallel convolutional layers, and each of them has a width of 8. Additionally, there are 101 layers in the network, which amount to a total of 88.8M parameters. We trained this model from scratch for 200 epochs, with early stopping enabled and a batch size of 8 samples, using a cross-entropy loss function. We used an Adam [10] optimizer with a learning rate of 0.001, which was reduced to 0.0001 after the 100th epoch. The same training recipe was used to train ResNext on all of the datasets in our experiments (Sec. 6).
5.2 Relative Object Detection
@@
As mentioned in Section 4.2, the goal is to predict a set of coordinates x and y for the chess pieces relative to the chessboard grid in the image. Thus, traditional object detection models that use Region Proposal Networks (RPNs) [21] or anchor boxes [20] are not suitable for this task, since they output absolute image coordinates. However, a single end-to-end object detection model, like Detection Transformer (DETR) [2], that directly predicts bounding boxes and class labels for objects in an image could be employed. DETR uses a transformer encoder-decoder architecture, with the encoder taking as input a feature map produced by a convolutional backbone network and the decoder generating the final predictions using self-attention mechanisms to attend to different parts of the feature map.
@@
For our experiments, we attempted to train a modified version of DETR that predicts relative object coordinates and omits the height and width dimensions for the bounding boxes of the traditional object detection task. ResNext101_32d was used as a backbone network for feature extraction. We set the number of queries (i.e. the maximum number of objects that DETR can detect in an image) to 32, since each chessboard can have at most 32 chess pieces on top of it. DETR also requires a separate class for “background”, which in our case corresponds to “empty” squares. Thus, the number of classes that the model is trained to predict is 13 (i.e. 12 piece types and background). We trained the model from scratch for a total of 800 epochs, with early stopping enabled and a batch size of 8 samples, using DETR’s default bipartite matching loss for set predictions, which takes into account both the class prediction and the similarity of the predicted and ground truth coordinates. We used an AdamW [13] optimizer with separate learning rates for the backbone network and the encoder-decoder architecture. In particular, the initial learning rates were set to $10^{-5}$ and $10^{-6}$ for the encoder-decoder and backbone, respectively, and a scheduler was used to reduce both by a factor of 10 every 300 epochs. Furthermore, gradient clipping was used with a threshold of $0.1$.
@@
However, the training of this modified DETR variant for chess recognition did not yield optimal results, with the model being unable to successfully detect chess pieces in the images of ChessReD. This issue could potentially be linked to DETR’s inherent limitation in detecting small objects [2, 32], especially when considering the intricacies of the dataset (e.g. occlusions) and the relatively small sizes of individual pieces. Due to the unsuccessful convergence of the DETR variant, it will not be used in the experiments of Section 6. Nevertheless, end-to-end relative object detection with transformers is a promising area that should be further investigated, with the focus being on refining the model architecture ([32]) or the training objective.
@@ 6.1 Exp1: Comparison with the state-of-the-art
To the best of our knowledge, the current state-of-the-art approach in chess recognition, namely Chesscog, was introduced in Wölflein and Arandjelović [28]. In their experiments, Chesscog achieved a 93.86% accuracy in chess recognition on a synthetic dataset [27] rendered in Blender[5], with a 0.23% per-square error rate. Additionally, the authors introduced a few-shot transfer learning approach to unseen chess sets and the system demonstrated a 88.89% accuracy and 0.17% per-square error rate, when tested on a set of previously unseen images of chessboards. In this section, we will compare the performance of our approach with that of Chesscog’s, both on their Blender dataset and on our newly introduced ChessReD.
@@ 6.1.1 Current SOTA: Chesscog
Chesscog attempts to solve the chess recognition task using a pipeline that involves chessboard detection, square localization, occupancy classification, and piece classification. It leverages the geometric nature of the chessboard to detect lines and employs a RANSAC-based algorithm to compute a projective transformation of the board onto a regular grid. Subsequently, individual squares are localized based on the intersection points and an occupancy classifier is used on each individual square. Finally, the pieces on the occupied squares are classified into one of 12 classes, using a pre-trained piece classifier. The piece classifier is used on image patches of the squares that are cropped based on a heuristic approach that extends the bounding boxes based on the square’s location on the chessboard. It is also important to mention that during inference the user must manually input the specific player’s perspective (i.e. “white” or “black”) to determine the orientation of the board.
6.1.2 How does the classification approach compare to Chesscog on the synthetic Blender dataset?
First, we evaluate the performance of our classification approach on the Blender dataset and compare it with that of Chesscog. The Blender dataset comprises a set of 4,888 synthetic chessboard images with distinct piece configurations, multiple lighting conditions, a limited range of viewing angles (between 45° and 60° to the board), and images taken only from the players’ perspectives. We trained our ResNeXt model following the recipe described in Section 5.1 on the dataset’s training samples. Subsequently, we evaluated our trained model’s performance on the test set. The first two columns of Table 3 demonstrate the evaluation results for both approaches on the Blender dataset. We use the same evaluation metrics as in Wölflein and Arandjelović [28].
@@ 6.1.2 How does the classification approach compare to Chesscog on the synthetic Blender dataset?
Mean incorrect squares per board
0.15 1.19 42.87 3.40 12.96 3.35
Boards with no mistakes (%)
93.86% 39.76% 2.30% 15.26% 6.69% 15.30%
Boards with $\leq 1$ mistake (%)
99.71% 65.20% 7.79% 25.92% 22.67% 27.04%
Per-square error rate (%)
0.23% 1.86% 73.64% 5.31% 39.57% 5.24%
@@ 6.1.3 How does the classification approach compare to Chesscog on the real ChessReD dataset?
In this section, we compare the performance of our approach with that of Chesscog’s on our ChessReD. We trained our ResNeXt model, using again the recipe of Section 5.1, and finetuned the Chesscog classifiers as mentioned in [28], using two images of the starting position from both players’ perspectives. Furthermore, for a fair comparison we needed to take into account that Chesscog cannot infer the orientation of the chessboard and requires for it to be manually inputted. Since this information is not available in our dataset, we address it by generating all possible orientations for the detected chessboards during evaluation.
@@
6.2 Exp2: Does the classification approach rely on chessboard marks for recognition?
(c) With marks
(d) Without marks
Figure 3: Sample pair of images for the ablation study
One significant advantage of our approaches is that they do not require any further input to determine the orientation of the chessboard in an image. Yet, the visual cues used by the models to deduce the chessboard’s orientation remain unclear. To this end, we conducted an ablation study to investigate whether the ResNeXt model relies on specific marks of the chessboard (e.g. bishop logo in Fig. 2(c)) to determine the board’s orientation and subsequently recognize the chess configuration.
@@ 7 Discussion
The evaluation on the Blender dataset revealed that Chesscog outperforms our classification approach. However, further experimentation on our newly introduced ChessReD showcased a shift in both methods’ performances, with ResNeXt surpassing Chesscog across all metrics. It is evident that Chesscog’s low chessboard detection rate (34.38%), which is attributed to the diverse angles and occlusions introduced by our dataset, significantly contributed to that shift, while the specific range of angles used in the Blender dataset, enabled Chesscog to successfully detect the chessboard in all cases and achieve a remarkable end-to-end performance.

### v3

@@ 4 METHOD: END-TO-END CHESS RECOGNITION
4.1 Implementation Details
Classification
@@ 4.1 Implementation Details
For the classification approach, we employed a ResNeXt [Xie et al., 2017] model. Their introduced concept of “cardinality” (i.e. the number of parallel branches used in each residual block) both enables deeper architectures with reduced computation complexity and allows complex representations to be learned by aggregating the information of the multiple branches. Because of these modifications, this family of models can achieve impressive results in image classification.
@@
For our experiments, we trained the resnext101_32d variant, which uses a cardinality of 32 and a width (i.e. number of filters) of 8. This means that each residual block in the network contains 32 parallel convolutional layers, and each of them has a width of 8. Additionally, there are 101 layers in the network, which amount to a total of 88.8M parameters. We trained this model from scratch for 200 epochs, with early stopping enabled and a batch size of 8 samples, using a cross-entropy loss function. We used an Adam [Kingma and Ba, 2014] optimizer with a learning rate of 0.001, which was reduced to 0.0001 after the 100th epoch. The same training recipe was used to train ResNext on all of the datasets in our experiments (Section 5).
Relative Object Detection
@@
As mentioned in Section 4, the goal is to predict a set of coordinates x and y for the chess pieces relative to the chessboard grid in the image. Thus, traditional object detection models that use Region Proposal Networks (RPNs) [Ren et al., 2015] or anchor boxes [Redmon et al., 2016] are not suitable for this task, since they output absolute image coordinates. However, a single end-to-end object detection model, like Detection Transformer (DETR) [Carion et al., 2020], that directly predicts bounding boxes and class labels for objects in an image could be employed. DETR uses a transformer encoder-decoder architecture, with the encoder taking as input a feature map produced by a convolutional backbone network and the decoder generating the final predictions using self-attention mechanisms to attend to different parts of the feature map.
@@
For our experiments, we attempted to train a modified version of DETR that predicts relative object coordinates and omits the height and width dimensions for the bounding boxes of the traditional object detection task. ResNext101_32d was used as a backbone network for feature extraction. We set the number of queries (i.e. the maximum number of objects that DETR can detect in an image) to 32, since each chessboard can have at most 32 chess pieces on top of it. DETR also requires a separate class for “background”, which in our case corresponds to “empty” squares. Thus, the number of classes that the model is trained to predict is 13 (i.e. 12 piece types and background). We trained the model from scratch for a total of 800 epochs, with early stopping enabled and a batch size of 8 samples, using DETR’s default bipartite matching loss for set predictions, which takes into account both the class prediction and the similarity of the predicted and ground truth coordinates. We used an AdamW [Loshchilov and Hutter, 2017] optimizer with separate learning rates for the backbone network and the encoder-decoder architecture. In particular, the initial learning rates were set to $10^{-5}$ and $10^{-6}$ for the encoder-decoder and backbone, respectively, and a scheduler was used to reduce both by a factor of 10 every 300 epochs. Furthermore, gradient clipping was used with a threshold of $0.1$.
@@
However, the training of this modified DETR variant for chess recognition did not yield optimal results, with the model being unable to successfully detect chess pieces in the images of ChessReD. This issue could potentially be linked to DETR’s inherent limitation in detecting small objects [Carion et al., 2020, Zhu et al., 2020], especially when considering the intricacies of the dataset (e.g. occlusions) and the relatively small sizes of individual pieces. Due to the unsuccessful convergence of the DETR variant, it will not be used in the experiments of Section 5. Nevertheless, end-to-end relative object detection with transformers is a promising area that should be further investigated, with the focus being on refining the model architecture ([Zhu et al., 2020]) or the training objective.
@@ 5.1 Exp1: Comparison with the State-of-the-art
To the best of our knowledge, the current state-of-the-art approach in chess recognition, namely Chesscog, was introduced in Wölflein and Arandjelović [Wölflein and Arandjelović, 2021]. In their experiments, Chesscog achieved a 93.86% accuracy in chess recognition on a synthetic dataset [Wölflein and Arandjelovic, 2021] rendered in Blender[Community, 2018], with a 0.23% per-square error rate. Additionally, the authors introduced a few-shot transfer learning approach to unseen chess sets and the system demonstrated a 88.89% accuracy and 0.17% per-square error rate, when tested on a set of previously unseen images of chessboards. In this section, we will compare the performance of our approach with that of Chesscog’s, both on their Blender dataset and on our newly introduced ChessReD.
@@ 5.1.1 Current SOTA: Chesscog
Chesscog [Wölflein and Arandjelović, 2021] does chess recognition using a pipeline that involves chessboard detection, square localization, occupancy classification, and piece classification. It uses the geometric nature of the chessboard to detect lines and employs a RANSAC-based algorithm to compute a projective transformation of the board onto a regular grid. Subsequently, individual squares are localized based on the intersection points and an occupancy classifier is used on each individual square. Finally, the pieces on the occupied squares are classified into one of 12 classes, using a pre-trained piece classifier. The piece classifier is used on image patches of the squares that are heuristically cropped by extending the bounding boxes based on the square’s location on the chessboard. During inference, the user must manually input the specific player’s perspective (i.e. “white” or “black”) to determine the orientation of the board.
5.1.2 Comparison to Chesscog on their Synthetic Blender Dataset
First, we compare and evaluate the performance of our classification approach on Chesscog’s synthetic Blender dataset [Wölflein and Arandjelović, 2021]. The Blender dataset comprises a set of 4,888 synthetic chessboard images with distinct piece configurations, multiple lighting conditions, a limited range of viewing angles (between 45° and 60° to the board), and images taken only from the players’ perspectives. We trained our ResNeXt model following the recipe described in Section 4.1 on the dataset’s training samples. Subsequently, we evaluated our trained model’s performance on the test set. The first two columns of Table 3 demonstrate the evaluation results for both approaches on the Blender dataset. We use the same evaluation metrics as in Wölflein and Arandjelović [Wölflein and Arandjelović, 2021].
@@ 5.1.2 Comparison to Chesscog on their Synthetic Blender Dataset
Mean incorrect squares per board 0.15 1.19 42.87 3.40 12.96 3.35
Boards with no mistakes (%) 93.86% 39.76% 2.30% 15.26% 6.69% 15.30%
Boards with $\leq 1$ mistake (%) 99.71% 65.20% 7.79% 25.92% 22.67% 27.04%
Per-square error rate (%) 0.23% 1.86% 73.64% 5.31% 39.57% 5.24%
@@ 5.1.3 How does the Classification Approach Compare to Chesscog on the Real ChessReD Dataset?
In this section, we compare the performance of our approach with that of Chesscog’s on our ChessReD. We trained our ResNeXt model, using again the recipe of Section 4.1, and finetuned the Chesscog classifiers as mentioned in [Wölflein and Arandjelović, 2021], using two images of the starting position from both players’ perspectives. Furthermore, for a fair comparison we needed to take into account that Chesscog cannot infer the orientation of the chessboard and requires for it to be manually inputted. Since this information is not available in our dataset, we address it by generating all possible orientations for the detected chessboards during evaluation.
@@
5.2 Ablation: Chessboard Markings
(a) With marks
(b) Without marks
Figure 4: Sample pair of images for the ablation study.
One significant advantage of our approaches is that they do not require any further input to determine the orientation of the chessboard in an image. Yet, the visual cues used by the models to deduce the chessboard’s orientation remain unclear. To this end, we conducted an ablation study to investigate whether the ResNeXt model relies on specific marks of the chessboard (e.g. bishop logo in Figure 3(a)) which can, for example be detected [Karaoglu et al., 2017], to subsequently determine the board’s orientation and recognize the chess configuration.
@@ 6 DISCUSSION
The evaluation on the Blender dataset revealed that Chesscog outperforms our classification approach. However, further experimentation on our newly introduced ChessReD showcased a shift in both methods’ performances, with ResNeXt surpassing Chesscog across all metrics. It is evident that Chesscog’s low chessboard detection rate (34.38%), which is attributed to the diverse angles and occlusions introduced by our dataset, significantly contributed to that shift, while the specific range of angles used in the Blender dataset enabled Chesscog to successfully detect the chessboard in all cases and achieve a remarkable end-to-end performance.
@@ 6.2 Conclusion
ACKNOWLEDGEMENTS
This project is supported in part by NWO (project VI.Vidi.192.100).
APPENDIX
Figure 5: Early-game (less than 30 moves) samples from ChessReD.
Figure 6: Mid-game (more than 30 and less than 75 moves) samples from ChessReD.
Figure 7: End-game (more than 75 moves) samples from ChessReD.
s9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">